# Hybrid Training for Enhanced Multi-task Generalization in Multi-agent Reinforcement Learning

### v1

@@ Abstract
In multi-agent reinforcement learning (MARL), achieving multi-task generalization to diverse agents and objectives presents significant challenges. Existing online MARL algorithms primarily focus on single-task performance, but their lack of multi-task generalization capabilities typically results in substantial computational waste and limited real-life applicability. Meanwhile, existing offline multi-task MARL approaches are heavily dependent on data quality, often resulting in poor performance on unseen tasks. In this paper, we introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. Specifically, our framework extracts potential general skills from offline multi-task datasets. We then train policies to select the optimal skills under the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interactions. We empirically demonstrate that our framework effectively extracts and refines general skills, yielding impressive generalization to unseen tasks. Comparative analyses on the StarCraft multi-agent challenge show that HyGen outperforms a wide range of existing solely online and offline methods.
Introduction
@@
Multi-agent reinforcement learning (MARL) has drawn broad attention for addressing problems in areas such as multi-robot systems [22, 53], video game AIs [35, 4], and autonomous driving [59]. Most existing MARL algorithms remain narrow, in that they focus on optimizing performance for specific tasks [30, 46], resulting in a significant gap between their poor multi-task generalization abilities and the variability of MARL tasks in real-world scenarios. Training specific agents from scratch for each task using MARL algorithms remains very costly and inefficient. Therefore, developing a generalized multi-task MARL algorithm is crucial to address these inefficiencies and improve scalability across diverse MARL tasks.
Two significant obstacles currently limit generalization in multi-task MARL. First, the restrictive model architectures in most MARL algorithms, characterized by fixed input and output dimensions of their neural architectures, fail to accommodate the variability of inputs and outputs across different tasks [20]. Recent online multi-task MARL works primarily focus on training across a predefined set of tasks simultaneously [34, 23] or on fine-tuning pre-trained policies for specific target tasks [20, 63, 36]. Although these approaches utilize a universal input network architecture to address the first obstacle and show promising performance on certain tasks, they fail to resolve another issue of significantly varying policies across different tasks. This results in their learned policies being limited to training tasks and unable to transfer knowledge from source to unseen tasks without further fine-tuning. Offline multi-task MARL [60] involves extracting skills from static datasets and training policies that select and reuse these skills in new tasks, underscoring the potential of leveraging generalizable skills from offline data. However, the effectiveness of these offline methods is often sensitive to the quality of their training dataset. Specifically, when the dataset lacks sufficient optimal or diverse trajectories, agents struggle to learn general skills and optimal policies for source tasks, limiting their performance and generalization capabilities in new tasks. Recent advancements in hybrid reinforcement learning (RL) [12, 19, 50] have shown that extracting skills or behaviors from offline data and then reusing them in online single-agent environments offers a potential solution to addressing issues existing in currently purely online or offline multi-task MARL approaches. However, applications of such frameworks in multi-agent systems remain rare.
@@
In this paper, we propose HyGen: Hybrid Training for Enhanced Multi-Task Generalization, a novel hybrid multi-task MARL approach combining both online environment interaction and offline datasets, as depicted in Figure 1. HyGen first extracts general skills from multi-task offline datasets, and then relies on hybrid training to learn generalizable policies for selecting optimal skills. These general skills and trained policies can then be applied to unseen tasks, as depicted in Figure 1. Specifically, HyGen first extracts general skills using a global trajectory encoder and actions encoders. The global trajectory encoder extracts a set of general skills common across different tasks from multi-task offline datasets, while the action decoders learn to delineate different agent actions with the discovered skills. We then train policies to select the optimal skills to maximize the global return via the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interaction experiences, refining the action decoders to make our skills unconstrained by prior data. Our proposed hybrid training paradigm is unique in how it integrates online interactions with offline data. Unlike RLPD [2], which uses a fixed ratio to sample data from both the online replay buffer and the offline data buffer, our method employs a linearly decreasing ratio. This strategy leverages the efficiency of offline learning initially and progressively incorporates the diversity of online interactions for exploration. Our refined skills during hybrid training are significant compared to previous works in offline multi-task multi-agent skill discovery [60], where skills are constrained to offline data. We finally present empirical results on the Starcraft Multi-Agent Challenge (SMAC), where we show that HyGen achieves remarkable generalization to unseen tasks by discovering general skills and learning high-quality policies, outperforming existing state-of-the-art online and offline multi-task MARL methods.
Related Works
Multi-task MARL
@@
Multi-task MARL methods are more adaptable and efficient than single-task MARL due to knowledge reuse [44] across various tasks. However, reusing knowledge across different tasks comes with its own set of challenges, e.g., varying input and output dimensions, which requires networks with flexible structures, such as self-attention mechanisms [20, 63, 60]. The MT-MARL approach described in [34] distills single-task policies into a unified policy that excels across multiple related tasks. REFIL [23] employs randomized entity-wise factorization for multi-task learning. However, these online methods require simultaneous training across a predefined set of tasks, incurring high costs of online interactions. UPDeT [20] leverages transformer-based value networks to accommodate changes in populations and inputs but requires additional online fine-tuning for new tasks. ODIS [60], an offline multi-task skill-based MARL method, also utilizes transformer-based networks. While this approach generalizes to unseen tasks without additional fine-tuning by reusing general skills, its performance is often constrained by the quality of the dataset. Achieving great generalization ability in unseen tasks remains a challenge.
Skill Discovery in MARL
@@
Skill discovery is an effective approach for tackling complex tasks due to its ability to identify and build a library of skills, often without relying on extrinsic rewards [10]. Recently, single-agent skill learning methods have been extended to MARL. Most skill-based MARL approaches [54, 15, 21, 29, 55] develop skills online to improve coordination. However, they do not emphasize reusing these skills for unseen tasks. ODIS [60] brings skill discovery to offline MARL, extracting generalizable skills from offline multi-task data. However, the skills discovered by ODIS are limited to the dataset they are derived from and often perform poorly on unseen tasks when the dataset quality is only moderate. Discovering high-quality reusable skills remains a significant challenge.
Hybrid Reinforcement Learning
@@
Hybrid RL [45] has been popular recently since it can take advantage of both purely online and offline methods. Recent efforts have focused on developing offline-to-online RL, a promising paradigm to reuse offline discovered skills [12] or offline learned behaviors [19, 61]. Other studies [26, 45, 32] have concentrated on adapting Q-learning to hybrid settings. Notably, research [26, 2] on integrating offline data and online interactions into a hybrid buffer aligns closely with our approach. The work in [26] introduces a balanced replay scheme that effectively utilizes online samples by leveraging relevant, near-on-policy offline samples. RLPD [2] employs symmetric sampling, where each batch comprises 50% data from the online replay buffer and 50% from the offline data buffer. However, the application of hybrid settings to multi-agent environments is still relatively unexplored.
Background
@@
Recent multi-task MARL works consider policy learning among two or several cooperative multi-agent tasks. In our settings, we focus on a multi-agent task set $\left\{\boldsymbol{\mathcal{T}}\right\}$ which contains tasks with varying team sizes. A multi-agent task $\mathcal{T}_{i}\in\left\{\boldsymbol{\mathcal{T}}\right\}$ can be described as a decentralized partially observable Markov decision process (Dec-POMDP) [31] consisting of a tuple $G=\left\langle I,S,A,P,\Omega,O,R,\gamma\right\rangle$. $i\in I\equiv\{1,\ldots,n\}$ is one of the agents and $s\in S$ describes the global state of the environment. At each time step, each agent $i\in I$ chooses an action $a^{i}\in A$, forming a joint action $\mathbf{a}\in\mathbf{A}\equiv A^{n}$. This causes a transition on the environment according to the state transition function $P\left(s^{\prime}\mid s,\mathbf{a}\right):S\times\mathbf{A}\times S\rightarrow[0,1]$. All agents would receive the reward according to the reward function $r(s,\mathbf{a}):S\times\mathbf{A}\rightarrow\mathbb{R}$ and $\gamma\in[0,1)$ is a discount factor. In a partially observable scenario, the agents could not get the global state $s$ but draw their individual observations $o$ $\in\Omega$ according to observation function $O(s,i):S\times I\rightarrow\Omega$. $\tau_{i}\in\boldsymbol{\tau}$ denotes the trajectory of agent $i$ which is an action-observation history $\left(o_{i}^{1},a_{i}^{1},\ldots,o_{i}^{t-1},a_{i}^{t-1},o_{i}^{t}\right)$.
We partition the multi-agent task set $\{\boldsymbol{\mathcal{T}}\}$ into source $\{\boldsymbol{\mathcal{T}}_{source}\}$ and target $\{\boldsymbol{\mathcal{T}}_{target}\}$ subsets. For source tasks, we combine their respective online interaction environments $G_{source}$ with offline datasets $\mathcal{D}$, comprising pre-collected agent trajectories $\boldsymbol{\tau}=(s,\mathbf{o},\mathbf{a},r,s^{\prime},\mathbf{o^{\prime}})$ to train our general agents. Post-training, these agents are directly deployed to execute target tasks in a zero-shot setting, without additional training or fine-tuning. Given the uncertain sources and variable quality of the offline datasets $\mathcal{D}$, they cannot be considered expert experiences.
Hybrid Training for Enhanced Multi-Task Generalization
@@
Unsupervised Offline General Skill Discovery
Good skills are expected to be general latent knowledge across different tasks in multi-agent reinforcement learning. We assume the skill $z_{i}$ for agent $i$ is a discrete variable from a finite skill set $\mathcal{Z}$, where the number of skills $|\mathcal{Z}|$ is a hyper-parameter. In this project, We use a pair of VAE-style [18] networks, which contain a multi-head attention global trajectory encoder and an action decoder to abstract the skills from multi-task offline datasets. In terms of basic functions, the global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$ extracts the information containing the global state $s$ and joint action $\mathbf{a}$ among each agent in the multi-task offline trajectories datasets $\mathcal{D}$ into general skill $z_{i}$ for agent $i$. There are different lengths of state $s$ and joint action $\mathbf{a}$ across different tasks. To handle this issue we reuse the task decompose module in UPDet [20]. After the decomposing, $(s,\mathbf{a})$ in different tasks is transferred into a group of entities $\mathbf{e}=\{e_{agent}^{1},e_{agent}^{2},\dots,e_{enemy}^{1},e_{enemy}^{2},\dots\}$ with same entity length. Because each entity $e^{i}$ only contains its own relevant environment information, we use $\mathcal{N}$-head self-attention to calculate the mutual influence between the entities containing all agents and the substitute entities. For each head $i$, we first compute separate query, key, and value projections: $Q_{i}=MLP_{Q}^{i}(\mathbf{e}),K_{i}=MLP_{K}^{i}(\mathbf{e}),V_{i}=MLP_{V}^{i}(\mathbf{e})$, then we can calculate the separate attention
$Attn_{i}(\mathbf{e})=\operatorname{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K_{i}}}}\right)V_{i},\quad d_{K_{i}}=\operatorname{dim}(K_{i})$
The final attention of whole entities is $Attn_{total}=\mathrm{concat}\{Attn_{1},Attn_{2},\dots,Attn_{n}\},n=\mathcal{N}$. Since self-attention computing does not change the relative position of entities in the group, we can extract the attention embedding for each agent to compute the general skill $z_{i}=MLP_{e}(Attn^{agent_{i}}),Attn^{agent_{i}}\in Attn_{total}$. It is noted that since each head in self-attention can learn to focus on different features of the input data and capture information from different representational subspaces [49], we can regard the latent information abstracted by each head as one skill. Therefore in this project, we set the number of skills $|\mathcal{Z}|$ equal to the number of attention heads $\mathcal{N}$.
@@
Following $\beta$-VAE [18], The training objective is to maximize the likelihood of the real action $a_{i}$ from data, along with the KL divergence [17] between $q(z_{i}|s,\mathbf{a},i)$ and a uniform prior $\tilde{p}\left(z_{i}\right)$ as a regularization. The regularization with a uniform distribution of cooperative strategies can prevent the state encoder from choosing similar skills for all inputs, thereby helping to discover distinguished skills. The final objective reads:
$L_{\mathrm{VAE}}\left(\theta_{s},\phi\right)=-\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\Bigg{[}$ $\sum_{i=1}^{n}\mathbb{E}_{q(z_{i}|s,\mathbf{a},i)}\Big{[}\log p\left(a_{i}\mid\tau_{i},z_{i}\right)\Big{]}$
$-\beta D_{\mathrm{KL}}\Big{(}q(\cdot)\|\tilde{p}(\cdot)\Big{)}\Bigg{]}$
where $\theta_{s}$ and $\phi$ denote the parameters of the global trajectory encoder and the action decoder respectively, and $\beta$ is the regularization coefficient. There is no need for any external rewards during the training process, so the first stage is unsupervised training. Figure 2 summarizes the skill discovery processing.
Hybrid High-level Policy Learning
After discovering general skills from multi-task offline datasets, we further learn general high-level policy to use these skills with hybrid training. In this work, we use a hybrid sampling approach, fully utilizing trajectory data generated by online environment exploration and existing trajectories in offline datasets during the data sampling process. Contrasting with [2, 12, 19, 50], who either entirely disregard offline data or blend offline and online data uniformly, our approach dynamically adjusts the proportion of data used during training. Given that the model parameters are initially near-random, leveraging offline trajectories predominantly at the outset provides a stable starting point for learning. As the model’s performance improves, it increasingly benefits from exploring the online environment, thereby gradually transitioning to a higher proportion of online data to refine policies. When the model performance is close to the performance bottleneck of offline trajectories, the model mainly relies on exploration in the online environment to obtain better policies. Online exploration trajectories account for the majority of the training batch, and offline trajectories maintain a small proportion. In practice, We implement a linear decay scheme for adjusting the hybrid ratio $R_{h}$, defined by the following equation:
$R_{h}=\max(R_{end},\frac{(R_{start}-R_{end})\cdot t}{N})$

### v2

@@ Abstract
In multi-agent reinforcement learning (MARL), achieving multi-task generalization to diverse agents and objectives presents significant challenges. Existing online MARL algorithms primarily focus on single-task performance, but their lack of multi-task generalization capabilities typically results in substantial computational waste and limited real-life applicability. Meanwhile, existing offline multi-task MARL approaches are heavily dependent on data quality, often resulting in poor performance on unseen tasks. In this paper, we introduce HyGen, a novel hybrid MARL framework, Hybrid Training for Enhanced Multi-Task Generalization, which integrates online and offline learning to ensure both multi-task generalization and training efficiency. Specifically, our framework extracts potential general skills from offline multi-task datasets. We then train policies to select the optimal skills under the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interactions. We empirically demonstrate that our framework effectively extracts and refines general skills, yielding impressive generalization to unseen tasks. Comparative analyses on the StarCraft multi-agent challenge show that HyGen obtains superior performances across a broad range of tasks. Code is available at https://github.com/Mr-Bright/HyGen.
1 Introduction
Multi-agent reinforcement learning (MARL) has drawn broad attention for addressing problems in areas such as multi-robot systems (huttenrauch2017guided, ; wang2022distributedreinforcementlearningrobot, ), video game AIs (peng2017multiagent, ; cao2012overview, ), and autonomous driving (yun2022cooperative, ). Most existing MARL algorithms remain narrow, in that they focus on optimizing performance for specific tasks (lowe2017multi, ; sunehag2017value, ), resulting in a significant gap between their poor multi-task generalization abilities and the variability of MARL tasks in real-world scenarios. Training specific agents from scratch for each task using MARL algorithms remains very costly and inefficient. Therefore, developing a generalized multi-task MARL algorithm is crucial to address these inefficiencies and improve scalability across diverse MARL tasks.
Two significant obstacles currently limit generalization in multi-task MARL. First, the restrictive model architectures in most MARL algorithms, characterized by fixed input and output dimensions of their neural architectures, fail to accommodate the variability of inputs and outputs across different tasks (hu2021updet, ). Recent online multi-task MARL works primarily focus on training across a predefined set of tasks simultaneously (omidshafiei2017deep, ; iqbal2021randomized, ) or on fine-tuning pre-trained policies for specific target tasks (hu2021updet, ; zhou2021cooperative, ; qin2022multi, ). Although these approaches utilize a universal input network architecture to address the first obstacle and show promising performance on certain tasks, they fail to resolve another issue of significantly varying policies across different tasks. This results in their learned policies being limited to training tasks and unable to transfer knowledge from source to unseen tasks without further fine-tuning. Offline multi-task MARL (zhang2022discovering, ) involves extracting skills from static datasets and training policies that select and reuse these skills in new tasks, underscoring the potential of leveraging generalizable skills from offline data. However, the effectiveness of these offline methods is often sensitive to the quality of their training dataset. Specifically, when the dataset lacks sufficient optimal or diverse trajectories, agents struggle to learn general skills and optimal policies for source tasks, limiting their performance and generalization capabilities in new tasks. Recent advancements in hybrid reinforcement learning (RL) (freed2023learning, ; hu2023unsupervised, ; wagenmaker2023leveraging, ) have shown that extracting skills or behaviors from offline data and then reusing them in online single-agent environments offers a potential solution to addressing issues existing in currently purely online or offline multi-task MARL approaches. However, applications of such frameworks in multi-agent systems remain rare.
@@ 1 Introduction
In this paper, we propose HyGen: Hybrid Training for Enhanced Multi-Task Generalization, a novel hybrid multi-task MARL approach combining both online environment interaction and offline datasets, as depicted in Figure 1. HyGen first extracts general skills from multi-task offline datasets, and then relies on hybrid training to learn generalizable policies for selecting optimal skills. These general skills and trained policies can then be applied to unseen tasks. Specifically, HyGen first extracts general skills using a global trajectory encoder and action encoders. The global trajectory encoder extracts a set of general skills common across different tasks from multi-task offline datasets, while the action decoders learn to delineate different agent actions with the discovered skills. We then train policies to select the optimal skills to maximize the global return via the centralized training and decentralized execution paradigm (CTDE). During this stage, we utilize a replay buffer that integrates both offline data and online interaction experiences, refining the action decoders to make our skills unconstrained by prior data. Our proposed hybrid training paradigm is unique in how it integrates online interactions with offline data. Unlike RLPD (ball2023efficient, ), which uses a fixed ratio to sample data from both the online replay buffer and the offline data buffer, our method employs a linearly decreasing ratio. This strategy initially leverages the efficiency of offline learning and progressively incorporates the diversity of online interactions for exploration. Our refined skills during hybrid training are significant compared to previous works in offline multi-task multi-agent skill discovery (zhang2022discovering, ; liu2025learning, ), where skills are constrained to offline data. We finally present empirical results on the Starcraft Multi-Agent Challenge (SMAC), where we show that HyGen achieves remarkable generalization to unseen tasks by discovering general skills and learning high-quality policies, outperforming existing state-of-the-art multi-task MARL methods.
2 Related Works
2.1 Multi-task MARL
@@
Multi-task MARL methods are more adaptable and efficient than single-task MARL due to knowledge reuse (silva2021transfer, ) across various tasks. However, reusing knowledge across different tasks comes with its own set of challenges, e.g., varying input and output dimensions, which requires networks with flexible structures, such as self-attention mechanisms (hu2021updet, ; zhou2021cooperative, ; zhang2022discovering, ). The MT-MARL approach described in (omidshafiei2017deep, ) distills single-task policies into a unified policy that excels across multiple related tasks. REFIL (iqbal2021randomized, ) employs randomized entity-wise factorization for multi-task learning. However, these online methods require simultaneous training across a predefined set of tasks, incurring high costs of online interactions. UPDeT (hu2021updet, ) leverages transformer-based value networks to accommodate changes in populations and inputs but requires additional online fine-tuning for new tasks. Offline multi-task skill-based MARL methods such as ODIS (zhang2022discovering, ) and HiSSD (liu2025learning, ), also utilize transformer-based networks. While these approaches can generalize to unseen tasks without additional fine-tuning by reusing skills, their performance is typically limited by the quality of the dataset. Achieving great generalization ability in unseen tasks remains a challenge.
2.2 Skill Discovery in MARL
@@
Skill discovery is an effective approach for tackling complex tasks due to its ability to identify and build a library of skills, often without relying on extrinsic rewards (eysenbach2018diversity, ). Recently, single-agent skill learning methods have been extended to MARL. Most skill-based MARL approaches (yang2019hierarchical, ; he2020skill, ; vmapd2022, ; liu2022heterogeneous, ; yang2024hierarchical, ) develop skills online to improve coordination. However, they do not emphasize reusing these skills for unseen tasks. ODIS (zhang2022discovering, ) and VO-MASD (chen2024variationalofflinemultiagentskill, ) extract transferable skills from offline multi-task data, whereas HiSSD (liu2025learning, ) jointly learns common and task-specific skills. However, the skills discovered by purely offline methods are limited to the dataset from which they are derived and often perform poorly on unseen tasks when the dataset quality is only moderate. Discovering high-quality reusable skills remains a significant challenge.
2.3 Hybrid Reinforcement Learning
@@
Hybrid RL (song2022hybrid, ) has been popular recently since it can take advantage of both purely online and offline methods. Recent efforts have focused on developing offline-to-online RL, a promising paradigm to reuse offline discovered skills (freed2023learning, ) or offline learned behaviors (hu2023unsupervised, ; zhang2023policyexpansionbridgingofflinetoonline, ). Other studies (DBLP:journals/corr/abs-2107-00591, ; song2022hybrid, ; niu2023trustsimulatordynamicsawarehybrid, ) have concentrated on adapting Q-learning to hybrid settings. Notably, research (DBLP:journals/corr/abs-2107-00591, ; ball2023efficient, ) on integrating offline data and online interactions into a hybrid buffer aligns closely with our approach. The work in (DBLP:journals/corr/abs-2107-00591, ) introduces a balanced replay scheme that effectively utilizes online samples by leveraging relevant, near-on-policy offline samples. RLPD (ball2023efficient, ) employs symmetric sampling, where each batch comprises 50% data from the online replay buffer and 50% from the offline data buffer. However, the application of hybrid settings to multi-agent environments is still relatively unexplored.
3 Background
@@
Recent multi-task MARL works consider policy learning among two or several cooperative multi-agent tasks. In our settings, we focus on a multi-agent task set $\left\{\boldsymbol{\mathcal{T}}\right\}$ which contains tasks with varying team sizes. A multi-agent task $\mathcal{T}_{i}\in\left\{\boldsymbol{\mathcal{T}}\right\}$ can be described as a decentralized partially observable Markov decision process (Dec-POMDP) (monahan1982state, ) consisting of a tuple $G=\left\langle I,S,A,P,\Omega,O,R,\gamma\right\rangle$. $i\in I\equiv\{1,\ldots,n\}$ is one of the agents and $s\in S$ describes the global state of the environment. At each time step, each agent $i\in I$ chooses an action $a^{i}\in A$, forming a joint action $\mathbf{a}\in\mathbf{A}\equiv A^{n}$. This causes a transition on the environment according to the state transition function $P\left(s^{\prime}\mid s,\mathbf{a}\right):S\times\mathbf{A}\times S\rightarrow[0,1]$. All agents would receive the reward according to the reward function $r(s,\mathbf{a}):S\times\mathbf{A}\rightarrow\mathbb{R}$ and $\gamma\in[0,1)$ is a discount factor. In a partially observable scenario, the agents could not get the global state $s$ but draw their individual observations $o$ $\in\Omega$ according to the observation function $O(s,i):S\times I\rightarrow\Omega$. $\tau_{i}\in\boldsymbol{\tau}$ denotes the trajectory of agent $i$ which is an action-observation history $\left(o_{i}^{1},a_{i}^{1},\ldots,o_{i}^{t-1},a_{i}^{t-1},o_{i}^{t}\right)$.
The task set $\{\boldsymbol{\mathcal{T}}\}$ is divided into subsets for source tasks $\{\boldsymbol{\mathcal{T}}_{source}\}$ and target tasks $\{\boldsymbol{\mathcal{T}}_{target}\}$. Source tasks involve a combination of online interaction environments $G_{source}$ and offline datasets $\mathcal{D}$, which consist of pre-collected agent trajectories $\boldsymbol{\tau}=(s,\mathbf{o},\mathbf{a},r,s^{\prime},\mathbf{o^{\prime}})$. These datasets are crucial for training agents to generalize across various scenarios. Upon successful training, these agents are then deployed directly to handle target tasks $\{\boldsymbol{\mathcal{T}}_{target}\}$ in a zero-shot setting, where they perform without any additional training or fine-tuning. This approach tests the agents’ ability to apply learned skills to new and potentially more complex tasks, evaluating their adaptability and generalization capabilities.
4 Hybrid Training for Enhanced Multi-Task Generalization
@@
4.1 Unsupervised Offline General Skill Discovery
@@
Useful skills are expected to be general latent knowledge across different tasks in MARL. We assume the skill $z_{i}$ for agent $i$ is a discrete variable from a finite skill set $\mathcal{Z}$, where the number of skills $|\mathcal{Z}|$ is a hyper-parameter. In this project, We use a pair of VAE-style (higgins2017beta, ) networks, which contain a multi-head attention global trajectory encoder and an action decoder to abstract the skills from multi-task offline datasets. In terms of basic functions, the global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$ extracts the information containing the global state $s$ and joint action $\mathbf{a}$ among each agent in the multi-task offline trajectories datasets $\mathcal{D}$ into general skill $z_{i}$ for agent $i$. There are different lengths of state $s$ and joint action $\mathbf{a}$ across different tasks. To handle this issue we reuse the task decompose module in UPDet (hu2021updet, ). After the decomposing, $(s,\mathbf{a})$ in different tasks is transferred into a group of entities $\mathbf{e}=\{e_{agent}^{1},e_{agent}^{2},\dots,e_{enemy}^{1},e_{enemy}^{2},\dots\}$ with same entity length. Because each entity $e^{i}$ only contains its own relevant environment information, we use $\mathcal{N}$-head self-attention to calculate the mutual influence between the entities containing all agents and the substitute entities. For each head $i$, we first compute separate query, key, and value projections: $Q_{i}=MLP_{Q}^{i}(\mathbf{e}),K_{i}=MLP_{K}^{i}(\mathbf{e}),V_{i}=MLP_{V}^{i}(\mathbf{e})$, then we can calculate the separate attention:
$Attn_{i}(\mathbf{e})=\operatorname{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K_{i}}}}\right)V_{i},\quad d_{K_{i}}=\operatorname{dim}(K_{i})$ (1)
The final attention of whole entities is $Attn_{total}=\mathrm{concat}\{Attn_{1},Attn_{2},\dots,Attn_{n}\},n=\mathcal{N}$. Since self-attention computing does not change the relative position of entities in the group, we can extract the attention embedding for each agent to compute the general skill $z_{i}=MLP_{e}(Attn^{agent_{i}}),Attn^{agent_{i}}\in Attn_{total}$. It is noted that since each head in self-attention can learn to focus on different features of the input data and capture information from different representational subspaces (vaswani2017attention, ), we can regard the latent information abstracted by each head as one skill. Therefore in this project, we set the number of skills $|\mathcal{Z}|$ equal to the number of attention heads $\mathcal{N}$.
@@
Following $\beta$-VAE (higgins2017beta, ), The training objective is to maximize the likelihood of the real action $a_{i}$ from data, along with the KL divergence (hershey2007approximating, ) between $q(z_{i}|s,\mathbf{a},i)$ and a uniform prior $\tilde{p}\left(z_{i}\right)$ as a regularization. The regularization with a uniform distribution of cooperative strategies can prevent the state encoder from choosing similar skills for all inputs, thereby helping to discover distinguished skills. The final objective reads:
[…]
where $\theta_{s}$ and $\phi$ denote the parameters of the global trajectory encoder and the action decoder respectively, and $\beta$ is the regularization coefficient. Figure 2 summarizes the skill discovery processing.
4.2 Hybrid High-level Policy Learning
Figure 3: Training framework during the high-level policy learning phase of HyGen. The hybrid buffer contains trajectories from the online buffer $\mathcal{R}$ and the offline dataset $\mathcal{D}$. The observation encoder extracts representations from local information. Meanwhile, the mixing network employs self-attention to accommodate varying input dimensions across different tasks.
After discovering general skills from multi-task offline datasets, we further learn a general high-level policy to use these skills with hybrid training. In this work, we use a hybrid sampling approach, fully utilizing trajectory data generated by online environment exploration and existing trajectories in offline datasets during the data sampling process. Contrasting with (ball2023efficient, ; freed2023learning, ; hu2023unsupervised, ; wagenmaker2023leveraging, ) which either entirely disregard offline data or blend offline and online data uniformly, our approach dynamically adjusts the proportion of data used during training. Given that the model parameters are initially near-random, leveraging offline trajectories predominantly at the outset provides a stable starting point for learning. As the model’s performance improves, it increasingly benefits from exploring the online environment, thereby gradually transitioning to a higher proportion of online data to refine policies. When the model performance is close to the performance bottleneck of offline trajectories, the model mainly relies on exploration in the online environment to obtain better policies. Online exploration trajectories account for the majority of the training batch, and offline trajectories maintain a small proportion. In practice, We implement a linear decay scheme for adjusting the hybrid ratio $R_{h}$, defined by the following equation:
$R_{h}=\max(R_{end},\frac{(R_{start}-R_{end})\cdot t}{N})$ (3)